OpenStudy (carniel):

1. Solve for x x^2+5x+16=-3x+4 2.Solve for x 2x^2+2x-11=3x-5 3.Factor 8x^3-y3 4.Factor 24x^3+81y^3 Please explain how you got answer and 1 question each person, so that everyone can get a medal for answering.

OpenStudy (anonymous):

1. X=-6, x=-2

OpenStudy (anonymous):

Your equation is: x^2+5x+16 = -3x+4 Simplified form is: x^2 + 8x + 12 = 0 The solutions of this equation are: X1 = -2 X2 = -6

OpenStudy (anonymous):

2. Your equation is: 2x^2+2x-11 = 3x-5 Simplified form is: 2x^2 - x - 6 = 0 The solutions of this equation are: X1 = 2 X2 = -3/2

OpenStudy (anonymous):

4. Result 24x3+81y3=3(2x+3y)(4x2−6xy+9y2) Explanation Step 1 : After factoring out 3 we have: 24x3+81y3=3(8x3+27y3) Step 2 : To factor 8x3+27y3 we can use sum of cubes formula: I3+II3=(I+II)(I2−I⋅II+II2) After putting I=2x and II=3y , we have: 8x3+27y3=(2x+3y)(4x2−6xy+9y2)

OpenStudy (anonymous):

3. Result 8x3−y3=(2x−y)(4x2+2xy+y2) Explanation To factor 8x3−y3 we can use difference of cubes formula: I3−II3=(I−II)(I2+I⋅II+II2) After putting I=2x and II=y , we have: 8x3−y3=(2x−y)(4x2+2xy+y2)

OpenStudy (carniel):

Thankyou :)

OpenStudy (carniel):

I thought I should only have 1 x